Parents wish to have $130,000 available for a child’s education. If the child is now 5 years old, how much money must be set aside at 3% compounded semiannually to meet their financial goal when the child is 18 ?
1
Expert's answer
2020-04-03T17:09:58-0400

"A=P(1+\frac{r}{100})^n"

here A is the amount after n intervals P is the principal amount r is the interest rate and n is the number of intervals

here number of years is 18-5=13 but as the rate is applied semiannualy so n =2x13 =26

let's substitute in the formula

"130000=P(1+\frac{3}{200})^{26}\\130000=P(1.4727)\\P=\frac{130000}{1.4727}=\$88273.24"